Fred Daley Co. applied the high-low method of cost estimation to professional labor hours billed to clients for the last 6 months of 2015. What is the estimated variable order-filling cost component per order? What is fixed cost? Professional Labor Hours Month Cost July August September October November December 3,560 $330,000 8,500 $577,000 7,510 $527,500 6,200 $462,000 9100$407,000 4,020 $353,000 What is the total cost for 4,800 customer orders?
Answer #1
Professional
Labor hours
Cost
High-level 8500 577000
Low-level 3560 330000
Difference 4940 247000
Estimated variable order-filling cost per order = 247000/4940 = $50
Fixed cost = 577000-(8500*50)= $152000
Total cost for 4800 customer order = 152000+(4800*50)= $392000